OpenStudy (anonymous):

Find the general term a.n for the geometric sequence a1 = -5 and a2 = 10

OpenStudy (anonymous):

@jim_thompson5910 can you help

jimthompson5910 (jim_thompson5910):

r = a2/a1 r = ??

OpenStudy (anonymous):

r = -2

jimthompson5910 (jim_thompson5910):

r = -2 good

jimthompson5910 (jim_thompson5910):

so your nth term is an = a1*r^(n-1) an = -5*(-2)^(n-1)

OpenStudy (anonymous):

would it be an = -5n(-2n + 2)?? @jim_thompson5910

OpenStudy (anonymous):

Sorry -5(-2n + 2)

jimthompson5910 (jim_thompson5910):

no you don't distribute, you just leave it as \[\Large a_{n} = -5(-2)^{n-1}\]

OpenStudy (anonymous):

thank you! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

So first find the common ratio. \[r = \frac{ a _{n} }{ a _{n-1} }\] \[r = \frac{ 10 }{ -5 }\] \[r = -2\] Now plug into the Formula. \[a _{n} = a _{1} \times r ^{n-1}\] \[a _{n} = -5 (-2) ^{n-1}\]

OpenStudy (anonymous):

thank you @some_someone